Outside temperature over a day can be modeled as a sinusoidal function. Suppose you know the high temperature of 88 degrees occurs at 3 PM and the average temperature for the day is 75 degrees. Find the temperature, to the nearest degree, at 6 AM. Preview degrees
Outside temperature over day and the average daily temperature first occurs at 8 AM. How many hours after midnight, to two decimal places, does the temperature first reach 58 degrees? a day can be modeled as a sinusoldal function. Suppose you know the temperature varies between 51 and 79 degrees during the Pretew hours
A Ferris wheel is 20 meters in diameter and boarded from a platform that is 1 meters above the ground. The slx o'clock position on the Ferris wheel is level with the loading platform. The wheel completes 1 full revolution in 4 minútes. How many minutes of the ride are spent higher than 16 meters above the ground? Preview minutes
Answer #1

given Tm=88o at 3 PM .Tavg=75o

The sinusoidal function can be modeled as

T(tToTsin(wt)

Where To is average temp in F.

we know that sin maximum occurs at 900.

But we have given Tmax=880F

from equation Tmax =T0+T=>88=75+T

therefore T=130F.

Now

wt T/2

t=15hr from 12.00am as reference time

w 2 15 30

on substituting the values

T(t7513sin ( 30

now for 6 pm t=12+6=18hr.

T18 T(18) 7513sin( 7512.364 87.364 30

Due to time constraint not possible answer at once please for remaining part re-post the rest.